OpenStudy (anonymous):

What is the average value of the function g(x) = 3^(cosx) on the closed interval [-pi, pi]

OpenStudy (anonymous):

\[1/(2\pi)\int\limits_{-\pi}^{\pi} 3^{\cos(x)} dx\]

OpenStudy (anonymous):

@sourwing Are we multiplying 3^cosxdx by 1/(2pi) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am not sure what the derivative of 3^(cosx) is. @sourwing

OpenStudy (anonymous):

Oh wait, 3^(cosx)dx = -3sin(x), right?

OpenStudy (anonymous):

you're integrating, not differenticating

OpenStudy (anonymous):

Okay, I got 1.325.